Find the measure of the angle described. The angle's measure is 26 degrees more than that of its complement? HELP!!!
Answer by macston(5194)

Angle plus it's complement=90
A=angle;A-26=complement
A+(A-26)=90 add 26 to each side
2A-26+26=90+26
2A=116 divide both sides by 2
A=58 ANSWER Angle is 58 degrees
Complement=A-26=58-26=32 degrees ANSWER Complement is 32 degrees
CHECK
Angle + Complement= 90 degrees
58 degrees+32 degrees=90 degrees
90 degrees=90 degrees